OpenStudy (anonymous):

método de Laplace y''-6y '9 y = ty (0) = 0 y' (0) = 1

OpenStudy (amistre64):

work the method, where are you stuck?

OpenStudy (amistre64):

also, fix your notation ...

OpenStudy (anonymous):

wait can u even speak english @marquito94

OpenStudy (amistre64):

Laplace works in other countries so i doubt english is needed

OpenStudy (anonymous):

@amistre64 true

OpenStudy (loser66):

@amistre64 but it is necessary when the Asker needs more explanation or makes question on the problem. For example, if I don't understand your question "Where are you stuck", how can I answer that question?

hartnn (hartnn):

gogle translate to the rescue!

OpenStudy (anonymous):

but he is is not speaking english he is speaking spanish i speak spanish

OpenStudy (loser66):

@hartnn a lot of fun when using google translate

OpenStudy (solomonzelman):

Yeah, google translate is off for so many things. I know this because I know other languages too.

OpenStudy (amistre64):

metodo seems more italian to me than spanish

OpenStudy (amistre64):

either way, the notation is off, something is missing in the equation.

hartnn (hartnn):

between 6y' and 9y whether it is + or -

hartnn (hartnn):

\( \Large y''-6y ' \pm 9 y = t \\ \Large y (0) = 0 , ~ y' (0) = 1\)

OpenStudy (amistre64):

\[L\{f''(t)\}=s^2L\{f'(t)\}-f'(0)-f(0)\] sounds familiar

OpenStudy (amistre64):

might be missing an s in there somplace lol

hartnn (hartnn):

\(\Large s^2 Y - s\times 0 -1 - 6sY +6(0) \pm 9Y = \dfrac{1}{s^2} \)

hartnn (hartnn):

yeah, its \(\Large s^2Y(s) -sy(0) -y'(0)\)

OpenStudy (amistre64):

:) thats the one ...

hartnn (hartnn):

\(\Large Y (s^2 -6y \pm 9) = \dfrac{1}{s^2} +1\)

hartnn (hartnn):

\(\Large y =L^{-1}[\dfrac{1+s^2}{(s^2+1)(s^2-6y \pm 9 )} ] \) i suspect it is +9 that would make it (s-3)^2

hartnn (hartnn):

\(\Large y =L^{-1}[\dfrac{1+s^2}{(s^2+1)(s^2-6y + 9 )} ]\) \(\Large y =L^{-1}[\dfrac{1+s^2}{(s^2+1)(s+3)^2} ]\) partial fractions

hartnn (hartnn):

****\(\Large y =L^{-1}[\dfrac{1+s^2}{(s^2+1)(s-3)^2} ]\)

hartnn (hartnn):

after partial fractions, its just looking up formulas in the laplace table